# Online Episodic Memory Visual Query Localization with Egocentric Streaming Object Memory

### v1

@@ 4.2 Object Memory Population (OMP)
The OT module is responsible for updating the localization of objects stored in $\mathcal{M}_{ego}$ at previous time steps, maintaining their unique IDs. At each frame $\mathbf{F}_{t}$, the OT module first retrieves all previously memorized objects $\mathbf{O}_{i}$ via $\mathcal{M}_{ego}.\texttt{read}\texttt{(}i\texttt{)}$, $\forall i$. The current bounding-box $\mathbf{b}_{i,t}$ is then computed by means of a visual object tracker [48, 42]. The visual tracker exploits the spatial and visual information stored in $\mathbf{O}_{i}$ to update the latest memorized bounding-box $\mathbf{b}_{i,\hat{t}},\hat{t}\leq t-1$ to $\mathbf{b}_{i,t}$. Alongside each $\mathbf{b}_{i,t}$, the tracker outputs confidence scores $s_{i,t}$ and only bounding-boxes with $s_{i,t}>\lambda_{ot}$ are retained. After filtering, the remaining bounding-boxes and their associated frame patches are written to memory as $\mathcal{M}_{ego}.\texttt{write}\left(\left[\mathbf{b}_{i,t},\phi_{i,t}\right]\right)$. If no objects are stored in $\mathcal{M}_{ego}$, i.e. when $t=0$ or until the new objects are inserted, the execution of the OT module is skipped. See the supplementary material for more details about the implementation of the OT module.
@@
The OD module is responsible for finding objects of interest in frame $\mathbf{F}_{t}$ which could be later queried by the user. OD is implemented as an object detector [45] which detects a set of $B$ bounding-boxes $\mathbf{b}_{k,t}$ of objects in the frame, coupled with confidence scores $s_{k,t}$, with $1\leq k\leq B$. After detecting the objects of interest, the OD module calls $\mathcal{M}_{ego}.\texttt{read}\texttt{(}t\texttt{)}$ to retrieve the bounding-boxes of all object instances already present in memory at time $t$. These are objects which may already have been tracked by the OT module at time $t$ from object instances at time $t-1$. The detected boxes are first filtered based on $s_{k,t}>\lambda_{od}$. Those that do not have an IoU $>\lambda_{iou}$ with any of the objects retrieved from the memory, are treated as localizations of new object instances. Their visual representations $\phi_{k,t}$ are extracted exploiting $\mathbf{b}_{k,t}$, and are then stored in memory by calling $\mathcal{M}_{ego}$.write($[\mathbf{b}_{k,t},\phi_{k,t}]$), while the others are discarded. See the supplementary material for more details.
@@ 4.3 Query Retrieval and Localization (QRL)
The Query (QRL)Object Retrieval and Query Localization algorithm is responsible for retrieving the response track $\mathbf{r}$ of a visual query $\mathbf{Q}$ within the memory $\mathcal{M}_{ego}$. This algorithm takes $\mathbf{Q}$ as input and outputs a sequence of contiguous bounding-boxes $\mathbf{b}_{i,t}$ for the object instance $\mathbf{O}_{i}$ in $\mathcal{M}_{ego}$ that visually matches $\mathbf{Q}$. The process begins by extracting feature representations $\Phi(\mathbf{Q})$ using a feature extraction function $\Phi(\cdot)$. Such representation is then compared against the features $\Phi(\phi_{i,t}),\phi_{i,t}\neq\emptyset$ of each object $\mathbf{O}_{i}$ in $\mathcal{M}_{ego}$ that are read via $\mathcal{M}_{ego}.\texttt{read}\texttt{(}i\texttt{)}$ with a feature similarity function $\Psi(\Phi(\mathbf{Q}),\Phi(\phi_{i,t}))\in[0,1]$. The comparisons generate a sequence of $[0,1]$ scores corresponding to each stored visual representation $\phi_{i,t}$. For each object $i$, these values are averaged into a single score $r_{i}$. If the maximum score $r_{i}$ exceeds the threshold $\lambda_{ret}$, the object $\mathbf{O}_{i}$ is considered a match for query $\mathbf{Q}$. Once matched, the latest sequence of contiguous bounding-boxes from $\mathbf{O}_{i}$’s spatio-temporal history is retrieved, returning it as the response track $\mathbf{r}$. Note that this process relies exclusively on $\mathcal{M}_{ego}$, and it does not require to access past frames from $\mathcal{V}_{ego}$. Figure 4 provides an illustration of the process performed by QRL.
@@ 5 Experimental Settings
Method Approach tAP $\uparrow$ tAP25 $\uparrow$ stAP $\uparrow$ stAP25 $\uparrow$ Succ $\uparrow$ Size $\downarrow$ Time $\downarrow$
1 SiamRCNN + KYS [18] offline 0.12 0.20 0.04 0.12 39.80 12.1 GB 60.3 m
2 STARK [48] offline - 0.10 - 0.04 18.70 12.1 GB 60.3 m
3 SiamRCNN [44] offline - 0.22 - 0.15 43.24 12.1 GB 5.1 m
4 CocoFormer [47] offline - 0.27 - 0.20 48.37 12.1 GB 60.3 m
5 VQLoC [22] offline - 0.31 - 0.22 55.89 12.1 GB 5.1 m
6 ESOM - ODoracle - OToracle online 0.68 0.73 0.65 0.68 81.92 7.1 MB 0.51 s
7 ESOM - ODyolo - OToracle online 0.14 0.21 0.12 0.19 40.55 11.2 MB 0.36 s
8 ESOM - ODoracle - OTegostark online 0.08 0.18 0.02 0.05 31.91 40.3 MB 0.49 s
9 ESOM - ODyolo - OTegostark online 0.0001 0.0003 0.0005 0.002 4.02 209.3 MB 3.1 s
Table 1: Comparison of ESOM for OEM-VQL versus offline VQL methods. Using ESOM for tackling OEM-VQL offers significant savings in memory storage and enhances the efficiency of query localization. Overall performance depends largely on the performance the OT and OD modules, where tracking accuracy has the largest impact.
@@
To evaluate ESOM, we used the Ego4D Episodic Memory VQ2D benchmark [18, 42], including $433$ hours of video from $54$ scenarios, featuring over 2200022000~{}2200022000 visual queries spanning 30003000~{}30003000 object categories. At present time, this is the only publicly available annotated dataset for Visual Query Localization. To train components, we also use EgoTracks [42] an egocentric video dataset sourced from Ego4D and aligned to the episodic memory benchmark, and EgoObjects [51], a large-scale egocentric dataset for object detection.
@@
[…]
@@
[…]
@@ 6 Results
6.1 State-of-the-art Comparison
OEM-VQL with ESOM vs offline VQL methods.
Table 1 compares different configurations of ESOM (rows 6-9) with state-of-the-art offline VQL methods (rows 1-5), highlighting significant advantages of ESOM in memory size and retrieval time. Offline methods require storing the entire video sequence, resulting in significant memory usage [^4] We consider a fixed memory of 12.1 GB to store a 5-minutes clip as required in the EGO4D benchmark. Longer clips would take more space.. ESOM reduces memory needs by up to 1000$\times$ when using oracular object discovery and tracking (row 6) and by 60$\times$ when using non-oracular models (row 9). Additionally, ESOM’s retrieval time decreases from the several minutes required by offline methods (5 m at least - row 5) to just seconds (3.1 s at most - row 9), enabling faster response times. This efficiency is essential for real-world applications with memory constraints or requirement of real-time performance.
Method tAP25 $\uparrow$ stAP25 $\uparrow$ Succ $\uparrow$ Size $\downarrow$ Time $\downarrow$
AMEGO [17] 0.0 0.0 0.67 241.7 MB 1.5 s
VideoAgent [11] 0.0002 0.001 3.03 1.2 GB 12.1 s
ESOM 0.0003 0.002 4.02 209.3 MB 3.1 s
Table 2: ESOM’s memory representation enables more accurate online visual query localizations. Comparison of ESOM with other methods to form object memories for OEM-VQL. (Best results are highlighted in bold).
Due to the higher complexity of EOM-VQL, as compared to standard offline VQL, results of offline and online methods are not directly comparable. Despite this fact, when ESOM is equipped with optimal discovery and tracking (row 6), we achieve significant results that even outperform offline methods (81.92 Succ. vs the best offline Succ. of 55.89 - row 4). While oracular settings are not indicative of real-world performance, these results show the great potential of performing VQL online with the proposed ESOM architecture. Replacing the oracle object discovery model (ODoracle) with the state-of-the-art detector ODyolo reduces accuracy by roughly 50%, from 81.92 to 40.55 (rows 7 vs 6). Note that, despite object detections being imperfect, these results are still competitive with most offline methods (e.g., SiamRCNN and CocoFormer in rows 3-4 achieve similar (s)tAP metrics and slightly higher Succ.), while bringing the aforementioned benefits in size and time. Replacing oracular tracking with OTegostark while keeping the optimal object detector ODoracle brings a decline of 60% (row 8 vs 6). Despite the reduction in performance, this combination still outperforms the offline STARK baseline (row 2). The combined use of non-oracular components, ODyolo and OTegostark, leads to a drop in performance of about 95%, emphasizing the limitations of current state-of-the-art object detection and tracking algorithms, in particular when they are employed to tackle a challenging downstream task such as the proposed OEM-VQL. Note that, while previous studies highlighted the challenges of addressing object detection [51] and visual object tracking [9, 10, 42] in egocentric vision, the de-facto standard for evaluating such models consists in assessing their performance in isolation [51, 42], which, we argue, may result in a distorted picture of the reliability of approaches for real-world applications. Besides setting a benchmark for the novel OEM-VQL task, our investigation also proposes a principled approach for evaluating object detection and tracking algorithms in egovision.
ESOM versus other approaches based on object memory representations.
Table 2 compares ESOM with two baselines obtained combining our QRL method with object memory representations recently proposed for tasks other than Online Episodic Memory. AMEGO [17] is designed to tackle interaction-based egocentric video understanding tasks and focuses on objects that the camera wearer has interacted with by hand, whereas VideoAgent [11] tackles video question answering and focuses on generic objects. The results show that, while AMEGO offers competitive memory usage and query localization times, its overall performance is very limited. This is likely due to its focus on interacted objects, which may result in discarding important future object queries. VideoAgent’s accuracy is closer to ESOM’s. However, this comes at a cost of much higher memory usage and query retrieval time, which ESOM reduces by approximately six and four times respectively. This is due to the focus of VideoAgent on generic objects, while ESOM is optimized to target object queries.
6.2 Analysis of ESOM and Ablations
OMP Configuration QRL tAP25 $\uparrow$ stAP25 $\uparrow$ Succ $\uparrow$ Time $\downarrow$
@@ 6.2 Analysis of ESOM and Ablations
ODoracle OToracle QRLsiamrcnn 0.73 0.68 81.92 0.51 s
QRLdino 0.64 0.57 76.00 0.56 s
ODyolo OToracle QRLsiamrcnn 0.21 0.19 40.55 0.36 s
QRLdino 0.20 0.12 35.53 0.4 s
ODoracle OTegostark QRLsiamrcnn 0.18 0.05 31.91 0.49 s
QRLdino 0.18 0.05 30.68 0.54 s
ODyolo OTegostark QRLsiamrcnn 0.0002 0.002 4.02 3.1 s
QRLdino 0.0002 0.002 3.79 3.4 s
Table 3: Effect of the QRL algorithm on performance. Performance of QRL methods across various OMP configuations. QRLsiamrcnn consistently achieves the highest accuracy, making it the best-performing model for query retrieval in OEM-VQ. (Best results are highlighted in bold).
@@
Table 3 compares the performance of the QRL algorithm under the two considered implementations. QRLsiamrcnn consistently performs best in matching a visual query with a memory of stored representations. Across all memory constructions produced by various OMP configurations, QRLsiamrcnn demonstrates superior query localization scores, emphasizing its strength and adaptability for the OEM-VQ task.
OMP Configuration $\phi_{i,t}$ tAP25 $\uparrow$ stAP25 $\uparrow$ Succ $\uparrow$ Size $\downarrow$
OD OT filter
ODoracle OToracle 0.73 0.68 81.92 7.1 MB
0.71 0.67 81.37 395 MB
ODyolo OToracle 0.21 0.19 40.55 11.2 MB
0.18 0.18 40.55 94.1 MB
ODoracle OTegostark 0.18 0.05 31.91 40.3 MB
0.17 0.05 29.78 350.2 MB
ODyolo OTegostark 0.0002 0.002 4.02 209.3 MB
0.0003 0.001 2.68 2.3 GB
Table 4: Memorizing good visual representations improves storage and retrieval. Keeping only good quality visual representation into the memory through filtering improves performance and reduces memory size. (Best results in bold).
Filtering visual representations helps storage and query retrieval.
We assessed how the filtering operation performed by means of the binary classifier during the .write() process affects memory size and the accuracy of QRL (Table 4). In the absence of the filtering operation, the memory is populated with the bounding boxes of all frames, along with the patches from the most recent response track, hence leading to an increased size. Downstream VLQ performance is generally improved by filtering, thanks to a more compact and uncluttered memory, with the added benefit of a decreased memory size.
OMP Config. AP $\uparrow$ Rec $\uparrow$ OMP Config. HOTA $\uparrow$ AssA $\uparrow$ tAP25 $\uparrow$ stAP25 $\uparrow$ Succ $\uparrow$ Size $\downarrow$ Time $\downarrow$
@@
OToracle 100.0 100.0 0.73 0.68 81.92 7.1 MB 0.51 s
ODoracle 100.0 100.0 OTmasa 50.1 25.2 0.02 0.01 9.38 191.7 MB 7.2 s
OTegostark 21.7 34.7 0.18 0.05 31.91 40.3 MB 0.49 s
OToracle 84.8 94.3 0.21 0.19 40.55 11.2 MB 0.36 s
ODyolo 13.6 30.7 OTmasa 6.6 11.6 0.001 0.001 3.75 46.7 MB 0.72 s
OTegostark 6.1 15.1 0.0003 0.002 4.02 209.3 MB 3.1 s
OToracle 87.5 94.3 0.15 0.12 33.33 14.7 MB 0.49 s
ODgdino 4.5 25.7 OTmasa 3.6 10.5 0.0 0.0 0.70 464.5 MB 2.5 s
OTegostark 5.7 15.1 0.0004 0.0 0.72 2.9 GB 24.1 s
@@
Table 5 presents the performance of ESOM under different configurations of the OD and OT modules, reporting module-specific and VQL task performance metrics: AP $\uparrow$ and Rec $\uparrow$ for object detection [^5] We merge all detections in a single “object” class., HOTA $\uparrow$ and AssA $\uparrow$ for object tracking [28], and (s)tAP25, Succ., Size and Time for downstream OEM-VQL.
The OD module has a noticeable impact on performance, as weaker detectors might miss important queries and add objects into memory that are unlikely to be queried later. Notably, the open-world ODgdino, which is designed to detect a wide range of object instances, exhibits low recall and AP for annotated queries, leading to weak retrieval results even when used with oracular tracking OToracle. In contrast, ODyolo, explicilty trained to target the queried objects enables effective memorization, resulting in more accurate QRL performance. Changes in the detector significantly affect memory size and processing time, with ODgdino, which detects a larger amount of objects, leading to a larger set of objects stored in memory. This expansion in memory size increases the time needed to localize a query.
OT modules have a more significant impact on performance. Indeed, changing the oracular tracker OToracle with OTegostark while maintaining the oracular object discovery ODoracle brings Succ. down from 81.92 to 31.91. Interestingly, downstream performance seems to be inversely correlated with general tracking performance, with OTmasa obtaining best HOTA $\uparrow$ scores (50.1 vs 21.7), but significantly smaller Succ. (9.38 vs 31.91), stAP25 (0.01 vs 0.05) and tAP25 (0.02 vs 0.18). Instead we observe correlation between VQL scores and association accuracy AssA, revealing that ID association is important for accurate response track prediction. Results with ODyolo and ODgdino for object discovery are more nuanced, but reveal similar inconsistencies between tracking and downstream metrics. Being explicitly optimized for egovision, OTegostark has the best downstream performance.
Figure 5: ESOM scales well while increasing memory size. Each row of plots shows how success score, storage space, and retrieval time change when performing QRL on memory built from progressively processed video segments. Results are shown for three different OMP configurations.

### v2

@@ 4.2 Object Memory Population (OMP)
The OT module is responsible for updating the localization of objects stored in $\mathcal{M}_{ego}$ at previous time steps, maintaining their unique IDs. At each frame $\mathbf{F}_{t}$, the OT module first retrieves all previously memorized objects $\mathbf{O}_{i}$ via $\mathcal{M}_{ego}.\texttt{read}\texttt{(}i\texttt{)}\>\forall i$, then estimates $\mathbf{b}_{i,t}$, the position of each object $i$ in frame $t$ by means of a visual object tracker [52, 46]. Alongside each $\mathbf{b}_{i,t}$, the tracker outputs confidence scores $s_{i,t}$ and only bounding boxes with $s_{i,t}>\lambda_{ot}$ are retained. After being filtered by the OT, the remaining bounding boxes and their associated frames are written to memory by calling $\mathcal{M}_{ego}.\texttt{write}\left(\left[t,\mathbf{b}_{i,t},\mathbf{F}_{i,t},c_{i,t}\right]\right)$. If no objects are stored in $\mathcal{M}_{ego}$, i.e. when $t=0$ or until the new objects are inserted, the execution of the OT module is skipped. See the supplementary material for more details about the implementation of the OT module.
@@
The OD module is responsible for finding objects of interest in frame $\mathbf{F}_{t}$ which could be later queried by the user. OD is implemented as an object detector [49] which detects a set of $B$ bounding boxes $\mathbf{b}_{k,t}$ of objects in the frame, coupled with confidence scores $s_{k,t}$, with $1\leq k\leq B$. After detecting the objects of interest, the OD module calls $\mathcal{M}_{ego}.\texttt{read}\texttt{(}t\texttt{)}$ to retrieve the bounding boxes of all object instances already present in memory at time $t$. These are objects which may already have been tracked by the OT module at time $t$ from object instances at time $t-1$. The detected boxes are first filtered based on $s_{k,t}>\lambda_{od}$. Those that do not have an IoU $>\lambda_{iou}$ with any of the objects retrieved from the memory, are treated as localizations of new object instances. Filtered $\mathbf{b}_{k,t}$ together with the associated frame $\mathbf{F}_{t}$, are then stored in memory by calling $\mathcal{M}_{ego}$.write($[t,\mathbf{b}_{k,t},\mathbf{F}_{t},c_{k,t}]$), while the others are discarded. See the supplementary material for details.
Retrieval Relevance.
To enhance the efficiency of the retrieval stage, each call to the .$\texttt{write}\texttt{()}$) method by OT and OD is preceded by a function that computes relevance scores $c_{i,t}$. These scores identify object patches $\phi_{i,t}$, extracted from $\mathbf{b}_{i,t}$ in $\mathbf{F}_{t}$, that contain sufficiently distinctive features for accurate instance recognition. We employ a binary classifier to assess whether each patch $\phi_{i,t}$ meets the quality threshold for effective retrieval (see Supp. material for details). If the classifier determines that a patch is suitable, $c_{i,t}=1$, otherwise, $c_{i,t}=0$.
@@ 4.3 Query Retrieval and Localization (QRL)
The Query Retrieval and Localization (QRL) algorithm is responsible for retrieving the visual response track $\mathbf{r}$ of a visual query $\mathbf{Q}$ within the memory $\mathcal{M}_{ego}$. This algorithm takes $\mathbf{Q}$ as input and outputs a sequence of contiguous bounding boxes $\mathbf{b}_{i,t}$ and frames $\mathbf{F}_{t}$ for the object instance $\mathbf{O}_{i}$ in $\mathcal{M}_{ego}$ that visually matches $\mathbf{Q}$. The process begins by extracting the query feature representation $\Phi(\mathbf{Q})$ using a feature extraction function $\Phi(\cdot)$. Such representation is hence compared against features $\Phi(\phi_{i,t})$ of object states $\mathbf{o}_{i,t}$ with $c_{i,t}=1$ using a feature similarity function $\Psi(\Phi(\mathbf{Q}),\Phi(\phi_{i,t}))\in[0,1]$. We average similarity scores related to multiple boxes $\mathbf{o}_{i,t}$ of the same object $i$ into a single score $r_{i}$. If the maximum score $r_{\overline{i}}$ across all objects exceeds the threshold $\lambda_{ret}$, the object $\mathbf{O}_{\overline{i}}$ is considered a match for query $\mathbf{Q}$. We hence return the most recent sequence of contiguous bounding boxes and frames from the spatio-temporal history of $\mathbf{O}_{\overline{i}}$ as the visual response track $\mathbf{r}$. Note that this process relies exclusively on $\mathcal{M}_{ego}$, and it does not require to access $\mathcal{V}_{ego}$. Figure 4 provides an illustration of the process performed by QRL.
@@ 5 Experimental Settings
Method OD OT %tAP25 $\uparrow$ %stAP25 $\uparrow$ %Succ $\uparrow$ Size $\downarrow$ Time $\downarrow$
1 AMEGO [21] ODhoi-det OTegostark 0.0 0.0 0.67 531.6 MB 1.5 s
2 VideoAgent [15] ODrt-detr OTbytetrack 0.02 0.1 3.03 13.7 GB 12.1 s
3 ESOM ODgdino OTmasa 0.002 0.001 0.70 1.3 GB 2.5 s
4 ESOM ODgdino OTegostark 0.04 0.002 0.72 8.8 GB 24.1 s
5 ESOM ODyolo OTmasa 0.1 0.1 3.75 383.6 MB 0.7 s
6 ESOM ODyolo OTegostark 0.03 0.2 4.02 1.7 GB 3.1 s
7 ESOM ODoracle OTegostark 18.3 5.1 31.91 591.4 MB 0.5 s
8 ESOM ODyolo OToracle 21.0 19.4 40.55 322.5 MB 0.6 s
9 ESOM ODoracle OToracle 73.2 68.1 81.92 507.4 MB 0.5 s
Table 1: Comparison of different configurations of ESOM for OVQ2D with respect to other online approaches. ESOM  outperforms competitors overall, with ODyolo-OTegostark being the best configuration according to success rate and ODyolo-OTmasa producing much faster inference and more compact memory sacrificing some performance. Grayed rows (7-9) report ESOM performance in oracle configurations. (Best results are highlighted in bold.)
@@
We evaluate ESOM on the Ego4D Episodic Memory VQ2D benchmark [22], including $433$ hours of video from $54$ scenarios, featuring $22K$ visual queries spanning $3K$ object categories. At present time, this is the only publicly available benchmark for Visual Query Localization. To train components, we also use EgoTracks [46], an egocentric video dataset sourced from Ego4D and aligned to the episodic memory benchmark, and EgoObjects [57], a large-scale egocentric object detection dataset.
@@
We tested different configurations of the proposed ESOM framework. In OMP, for the OT component, we used EgoSTARK [46], a STARK instance [52] fine-tuned on EgoTracks, as well as the MASA tracker [27]. To evaluate optimal tracking performance, we also considered ground-truth annotations from EgoTracks. We refer to these configurations as OTegostark, OTmasa, OToracle. For the OD module, we considered the off-the-shelf open-world detector GroundingDino [30], an instance of YOLOv10 [49] trained on curated versions of EgoTracks [46] and EgoObjects [57], and ground truth object detections from EgoTracks. These are denoted as ODgdino, ODyolo, ODoracle. In QRL, we implemented feature extraction $\Phi(\cdot)$ and similarity $\Psi(\cdot)$ using two configurations: Siam-RCNN with a classification head [48] or DINOv2 with cosine similarity [21, 34], referred to as QRLsiamrcnn and QRLdino. We obtained best results with OTegostark, ODyolo and QRLsiamrcnn, which identify the default configuration of ESOM where not otherwise specified. See the supplementary material for more implementation details.
@@
Each video clip $\mathcal{V}_{ego}$ in the evaluation dataset is fed to the OMP algorithm to construct the object memory $\mathcal{M}_{ego}$. Each visual query $\mathbf{Q}$ provided with the clip is then processed using the QRL algorithm, producing a visual response track $\mathbf{r}$, which we subsequently compare to the annotated ground-truth response track of bounding-boxes to compute query localization accuracy. We use the standard VQL evaluation metrics [22]: %tAP25 $\uparrow$, measures temporal alignment with the ground-truth at a 0.25 IoU threshold; %stAP25 $\uparrow$, assesses spatio-temporal precision a 0.25 spatio-temporal IoU threshold; Success (%Succ $\uparrow$) calculates the proportion of predictions with a minimum 0.05 spatio-temporal IoU with the ground-truth response track. All metrics are reported in the range [0, 100]. In addition to these standard measures, we evaluate efficiency in terms of storage size (Size $\downarrow$) in multiples of bytes (MB, GB) needed to store information that enables query localization, and retrieval time (Time $\downarrow$) in seconds (s) or minutes (m) that is required to localize a visual query.
@@ 6 Results
6.1 Online Visual Query 2D (OVQ2D) Benchmark
Table 1 benchmarks different configurations of ESOM (3-6) against two baselines (1-2) that apply our QRL method on object memory representations recently proposed for tasks other than Online Episodic Memory. AMEGO [21] is designed for interaction-based egocentric video understanding and focuses on objects that the camera wearer has interacted with by hand, whereas VideoAgent [15] tackles video question answering and focuses on generic objects. Since AMEGO and VideoAgent were not originally designed for the OVQ2D task, their baselines in Table 1 are adapted by integrating their respective object detection and tracking components—HOI-detector [43] and EgoStark for AMEGO, RT-DETR [56] and ByteTrack [54] for VideoAgent—with our memory and retrieval modules. Alongside real-world configurations, in (7-9), we also report baselines using oracular detection (7), object tracking (8) and both (9). Best results are obtained by the proposed ESOM when state-of-the-art detector (ODyolo) and tracker (OTegostark) are considered. Besides being competitive in terms of query retrieval accuracy (see columns 4-6), this configuration brings a fast retrieval of 3.1 s and a compact memory of 1.7 GB. Among competitors, VideoAgent achieves a lower success rate (3.03 vs 4.02) with much larger memory 13.7 vs 1.7 GB and slower retrieval time (12.5 vs 3.1 s), while AMEGO, focusing only on hand-interacted objects, achieves a very low success rate of 0.67. The choice of detector and tracker greatly affects the performance of ESOM. Indeed, using the generic object detector ODgdino, rather than our optimized ODyolo leads to very limited success rates, regardless of the tracker. OTmasa and OTegostark also lead to different results, with OTegostark bringing the best performance, while OTmasa allowing to obtain faster retrieval and a more compact memory sacrificing some retrieval accuracy. It is worth noting that the OVQ2D benchmark is very challenging for current approaches, with the best methods achieving a success rate of just 4.02 and very low object localization scores overall. The oracular results in rows 7-9 show that this is mainly due to the limited effectiveness of current object detection and tracking approaches in real-world egocentric videos. Indeed, pairing OTegostark with ground truth detections (row 7) brings to a success rate of 31.91, pairing a ODyolo object detection with an oracular tracker leads to a success rate of 40.55, while using both oracular detection and tracking leads to 81.92. Overall, these results highlight the potential of ESOM to tackle OVQ2D, but also highlight that efforts should be devoted to improve object detection and tracking. While prior studies highlight challenges in egocentric object detection [57] and tracking [11, 12, 46, 13], algorithms are typically evaluated in isolation [57, 46], which we argue may misrepresent their real-world reliability. Beyond OVQ2D, this benchmark, including its oracular baselines, can serve as a principled testbed to support the development and evaluation of object detection and tracking in real-world, downstream settings, rather than in isolation.
Method %Succ $\uparrow$ Size $\downarrow$ Time $\downarrow$
Offline Methods (VQ2D)
SiamRCNN + KYS [22] 39.80 12.1 GB 8.3 m
STARK [52] 18.70 12.1 GB 45 s
SiamRCNN [48] 43.24 12.1 GB 8.3 m
CocoFormer [51] 48.37 12.1 GB 8.3 m
VQLoC [26] 55.89 12.1 GB 41 s
PRVQL [14] 57.93 12.1 GB 50 s
ESOM (OVQ2D)
ESOM - ODoracle - OToracle 81.92 507.4 MB 0.5 s
Table 2: Comparison of ESOM in OVQ2D versus offline VQ2D methods. Note that VQ2D is much less challenging than OVQ2D, hence success rate is not directly comparable and these results are reported to discuss the differences between the two approaches. ESOM offers significant savings in memory storage and enhances the efficiency of query localization.
6.2 Comparison with Offline VQ2D
Table 2 shows state-of-the-art offline VQ2D methods. It is important to note that offline VQ2D and our online formulation are not directly comparable, OVQ2D is a much harder task in which methods do not know have access to the queried object in advance; hence these results are reported mainly to discuss the differences between the two approaches. Current offline methods require storing the entire video sequence, resulting in significant memory usageWe consider a fixed memory of 12.1 GB to store a 5-minutes clip as required in the EGO4D benchmark. Longer clips would take more space.. ESOM reduces memory needs by up 7$\times$ (1.7 vs 12.1GB) and increases inference speed by 13$\times$ when compared to VQLoC (3.1 vs 41 s). Memory and computational savings are even more significant when considering the oracular configuration, reported for reference, with 24$\times$ less memory and 80$\times$ increased inference speed. Interestingly, ESOM is competitive even in terms of success rate in oracular settings (81.92 vs 55.89 of VQLoC), which highlights the potential of OVQ2D and ESOM when tracking and detection are solved.
6.3 Detailed Analysis of ESOM and Ablations
OMP Configuration QRL %tAP25 $\uparrow$ %stAP25 $\uparrow$ %Succ $\uparrow$ Time $\downarrow$
@@ 6.3 Detailed Analysis of ESOM and Ablations
ODoracle OToracle QRLsiamrcnn 73.2 68.1 81.92 0.5 s
QRLdino 64.4 57.7 76.00 0.6 s
ODyolo OToracle QRLsiamrcnn 21.0 19.4 40.55 0.4 s
QRLdino 20.6 12.1 35.53 0.4 s
ODoracle OTegostark QRLsiamrcnn 18.3 5.1 31.91 0.5 s
QRLdino 18.3 5.1 30.68 0.5 s
ODyolo OTegostark QRLsiamrcnn 0.02 0.2 4.02 3.1 s
QRLdino 0.02 0.2 3.79 3.4 s
Table 3: Effect of the QRL algorithm on performance. Performance of QRL methods across various OMP configuations. QRLsiamrcnn consistently achieves the highest accuracy, making it the best-performing model for query retrieval in OVQ2D. (Best results are highlighted in bold).
@@
Table 3 compares the performance of the QRL algorithm under the two considered implementations. QRLsiamrcnn consistently performs best at matching a visual query with a memory of stored representations. Across all memory constructions produced by various OMP configurations, QRLsiamrcnn demonstrates superior query localization scores, emphasizing its strength and adaptability for the OVQ2D task.
Filtering visual representations helps retrieval accuracy and time.
We assessed how the computation of the relevance scores $c_{i,t}$ performed by means of the binary classifier before the .write() process affects the accuracy and speed of QRL (Table 4). In the absence of the retrieval relevance computation, all object representations stored in memory are considered during retrieval, increasing time as more similarity values need to be computed. Downstream OVQ2D performance is also improved by filtering, thanks to the exclusion of suboptimal object representations.
OMP Configuration Retr. %tAP25 $\uparrow$ %stAP25 $\uparrow$ %Succ $\uparrow$ Time $\downarrow$
OD OT Rel.
ODoracle OToracle 73.2 68.1 81.92 0.5 s
71.5 67.4 81.37 4.6 s
ODyolo OToracle 21.0 19.4 40.55 0.4 s
18.3 18.1 40.55 5.3 s
ODoracle OTegostark 18.3 5.1 31.91 0.5 s
17.4 5.0 29.78 24.2 s
ODyolo OTegostark 0.02 0.2 4.02 3.1 s
0.03 0.1 2.68 24.8 s
Table 4: Ablating the use of the retrieval relevance over different OMP configurations. Using relevant visual representations for retrieval (Retr. Rel.) improves time and accuracy of QRL. (Best results in bold).
OMP Config. AP $\uparrow$ Rec $\uparrow$ OMP Config. HOTA $\uparrow$ AssA $\uparrow$ %tAP25 $\uparrow$ %stAP25 $\uparrow$ %Succ $\uparrow$ Size $\downarrow$ Time $\downarrow$
@@
OToracle 100.0 100.0 73.2 68.1 81.92 507.4 MB 0.5 s
ODoracle 100.0 100.0 OTmasa 50.1 25.2 2.8 1.6 9.38 1.3 GB 7.2 s
OTegostark 21.7 34.7 18.3 5.1 31.91 591.4 MB 0.5 s
OToracle 84.8 94.3 21.0 19.4 40.55 322.5 MB 0.4 s
ODyolo 13.6 30.7 OTmasa 6.6 11.6 0.1 0.1 3.75 383.6 MB 0.7 s
OTegostark 6.1 15.1 0.03 0.2 4.02 1.7 GB 3.1 s
OToracle 87.5 94.3 15.7 12.2 33.33 355.1 MB 0.5 s
ODgdino 4.5 25.7 OTmasa 3.6 10.5 0.0 0.0 0.70 1.3 GB 2.5 s
OTegostark 5.7 15.1 0.04 0.0 0.72 8.7 GB 24.1 s
@@
Table 5 presents the performance of ESOM under different configurations of the OD and OT modules, reporting module-specific and VQL task performance metrics: AP $\uparrow$ and Rec $\uparrow$ for object detectionWe merge all detections in a single “object” class., HOTA $\uparrow$ and AssA $\uparrow$ for object tracking [32], and %(s)tAP25, Succ., Size and Time for downstream OVQ2D.
The OD module has a noticeable impact on performance, as weaker detectors might miss important queries and add objects into memory that are unlikely to be queried later. ODgdino, designed for open-world detection, has low recall and AP for annotated queries, leading to weak retrieval even with OToracle. In contrast, ODyolo, trained to detect queried objects, enables better memorization and improves QRL accuracy. Changes in the detector impact memory size and processing time; ODgdino detects more objects, expanding memory and increasing query localization time.
OT modules have a more significant impact on performance. Changing the oracular tracker OToracle with OTegostark while maintaining the oracular object discovery ODoracle brings Succ. down from 81.92 to 31.91. We observe an inverse correlation between downstream performance and general tracking performance. For instance, OTmasa achieves the highest HOTA score (50.1 vs. 21.7), yet yields significantly lower Succ. (9.38 vs. 31.91). This apparent discrepancy can be explained by the nature of the HOTA metric, which is defined as the geometric mean of DetA (detection accuracy within individual frames) and AssA (association accuracy across time) [32]. In our evaluations, DetA tends to be consistently lower than AssA — for example, ODoracle–OTegostark shows 13.7 vs. 34.7, and ODyolo–OTegostark yields 2.5 vs. 15.1—ultimately reducing overall HOTA scores. OVQ2D requires higher AssA than DetA, as accurate object association is essential for constructing coherent and temporally consistent response tracks in $\mathcal{M}_{ego}$. In cases where associations are incorrect, $\mathcal{M}_{ego}$ accumulates fragmented or duplicated tracklets, and DetA penalizes such false positives. ODoracle–OTmasa stands out as an exception, relying on ground truth detections and thus isolating the role of association, which explains its strong HOTA despite limited downstream success. Being explicitly optimized for egovision, OTegostark leads to the best downstream performance.
Figure 5: ESOM scales well while increasing memory size. Plots show how success score, storage space, and retrieval time change when building memory from progressively processed video segments. Results for three different OMP configurations.